how many grams of beryllium chloride would you need to add to 125 mL of water to make a 0.050 molal solution
1
Expert's answer
2019-04-08T07:55:13-0400

CM=n/V n=m/M M (BeCl2) = 80 g/mol

n (BeCl2) = 0.050 x 0.125 = 0.006 mol

m (BeCl2) = 0.006 x 80 = 0.5 g